# A Genetic Algorithm Approach to Anti-Jamming UAV Swarm Behavior thanks: This work was supported by national funds through FCT, Fundação para a Ciência e a Tecnologia, under project UIDB/50021/2020

### v1

@@ Abstract
In recent years, Unmanned Aerial Vehicles (UAVs) have brought a new true revolution to military tactics. While UAVs already constitute an advantage when operating alone, multi-UAV swarms expand the available possibilities, allowing the UAVs to collaborate and support each other as a team to carry out a given task. This entails the capability to exchange information related with situation awareness and action coordination by means of a suitable wireless communication technology. In such scenario, the adversary is expected to disrupt communications by jamming the communication channel. The latter becomes the Achilles heel of the swarm. While anti-jamming techniques constitute a well covered topic in the literature, the use of intelligent swarm behaviors to leverage those techniques is still an open research issue.
This paper explores the use of Genetic Algorithms (GAs) to jointly optimize UAV swarm formation, beam-steering antennas and traffic routing in order to mitigate the effect of jamming in the main coordination channel, under the assumption that a more robust and low data rate channel is used for formation management signaling. Simulation results show the effectiveness of proposed approach. However, the significant computational cost paves the way for further research.
@@ I Introduction
    The definition of a new optimization problem, which seeks to jointly optimizing the swarm formation, smart antenna configuration, and network traffic routing, to achieve more effective anti-jamming performance.
@@
    A Genetic Algorithm (GA) approach for optimizing UAV positions and antenna beam directions to maximize communication capacity in the presence of interference and/or malicious jamming.
@@ III System Model
    The positio of each UAV is defined in 2D Cartesian coordinates (x,y), and its smart antenna beam direction is represented as an angle between 0° and 360°;
@@
    Main channel beam-steering and formation coordination messages are conveyed through a dedicated omnidirectional low data rate robust signaling channel, which is resilient to jamming;
@@
A system model with single communication channel, where the signaling messages are also affected by jamming is relegated for future work.
@@ V Simulation Results
The simulation software was implemented in Python using various tools and libraries.
OpenAI Gym is an open-source library providing a standardized environment for developing and testing RL algorithms. It allows direct comparison of different algorithms in a consistent environment by handling critical simulation tasks, such as calculating node distances and antenna angles, implementing radiation diagrams for directional antennas, determining antenna gain, calculating reception and interference power, and assessing communication capacity between nodes.
Stable Baselines3 (SB3) is a library offering reliable implementations of RL algorithms in PyTorch. Its integration with OpenAI Gym enables direct comparisons between algorithms, and its customizable framework allows adaptation for specific research needs.
Seaborn, a Python data visualization library, provides a high-level interface for exploring complex datasets. This library enables effective visualization, interpretation, and comparison of the results obtained in the research.
@@ V-A Scenario 1: Static Positions with Beam-Steering Optimization
TABLE III: Study of GA parameterisation
@@
Figure 7 shows the evolution of the OF value for a beam-steering configuration, as well as the constant baseline value corresponding to the use of omnidirectional communication.
@@
It can be seen that that the use of beam-steering effectively improves communication performance. An example result is shown in Figure 8, which was obtained after a computation time of 29.3 s. This result corresponds to an average data rate of 2856 bps and a bottleneck data rate of 129 bps. In contrast, for the omnidirectional configuration, the average data rate was 0.89 bps and the bottleneck data rate was 0.12 bps.
Figure 8 shows the direction of the antenna taken by each UAV. It is easy to notice that the beam-steering solution favors pointing the beams away from the jammer (e.g., UAV0 and UAV3), and/or towards other UAVs (e.g., UAV1 and UAV2).
@@ V-B Scenario 2: Fixed Swarm Deployment Area with Beam-Steering and UAV Position Optimization
Figure 10: Scenario 2: convergence of the OF in the acquisition of positions.
@@
In the omnidirectional communication configuration, GA can obtain a very fine-tuned arrangement of the UAVs in a significantly shorter time interval. Its results are shown in Figure 12. Although the positioning of the UAVs using the beam-steering antenna configuration are obviously not ideal when compared to the omnidirectional configuration, there is a drastic decrease in communication performance when using the latter.
@@ V-C Scenario 3: Mobile Swarm Deployment Area
In this dynamic scenario, the UAV swarm continuously moves in a straight line towards some mission objective, while the UAVs can change their relative positions within its deployment area, as well adapt the beam-steering configuration accordingly. Figure 13 illustrates the scenario.
@@
Additional simulation parameters for Scenario 3 are listed in Table V.
@@
The initial UAV swarm situation is depicted in Figure 15. Selected snapshots of the UAV swarm movement are depicted in Figure 16. Despite processing time constraints, the GA could effectively cluster UAVs and direct antennas to mitigate jamming effects. Over the simulation, the sequence of objective function curves (omitted due to page limits) reflect a steady optimization of communication quality, as seen in evolving UAV topologies as the swarm operational area slides horizontally. The success in aligning antennas and adjusting positions demonstrates the GA capability to provide robust anti-jamming strategies within the given constraints, even if the found solutions were still not optimal. It must be noted that the GA seeks optimal UAV positions and beam-steering solutions for specific time instants of the swarm movement. Although the GA could improve its decisions for longer time windows, the latter would also mean that the swarm would adapt its topology and beam-steering less frequently. For very distant jammers, enlarging the time window is possible, since angular directions change slowly, otherwise, enlarging the time window will also lead to non-optimality. Optimization of the time window length is left for future work.

### v2

# Anti-Jamming based on Beam-Steering Antennas and Intelligent UAV Swarm Behavior

@@ Abstract
In recent years, Unmanned Aerial Vehicles (UAVs) have brought a new true revolution to military tactics. While UAVs already constitute an advantage when operating alone, multi-UAV swarms expand the available possibilities, allowing the UAVs to collaborate and support each other as a team to carry out a given task. This entails the capability to exchange information related with situation awareness and action coordination by means of a suitable wireless communication technology. In such scenario, the adversary is expected to disrupt communications by jamming the communication channel. The latter becomes the Achilles heel of the swarm. While anti-jamming techniques constitute a well covered topic in the literature, the use of intelligent swarm behaviors to leverage those techniques is still an open research issue. This paper explores the use of beam-steering antennas and UAV swarm formation adaptation in order to mitigate the effect of jamming in the main coordination channel, under the assumption that a more robust and low data rate channel is used for formation management signaling. A Genetic Algorithm (GA) approach is used to jointly optimize beam directions, swarm formation and traffic routing. Simulation results show the effectiveness of proposed approach. However, the significant computational cost paves the way for future research.
@@ I Introduction
    The definition of a novel optimization problem, which seeks to jointly optimizing the swarm formation, smart antenna configuration, and network traffic routing, to achieve more effective anti-jamming performance.
@@
    A Genetic Algorithm (GA) approach to show how optimizing UAV positions and antenna beam directions can maximize communication capacity in the presence of interference and/or malicious jamming.
@@ III System Model
    The position of each UAV is defined in 2D Cartesian coordinates (x,y), and its smart antenna beam direction is represented as an angle between 0° and 360°;
@@
    Main channel beam-steering and formation coordination messages are idealistically conveyed through a dedicated omnidirectional low data rate robust signaling channel, which is resilient to jamming;
@@
A system model with single communication channel, where the signaling messages are also affected by jamming is relegated for future work. The main objective of this paper is to show how swarm member behavior can mitigate jamming when properly coordinated.
@@ V Simulation Results
The simulation software was implemented in Python. Seaborn, a Python data visualization library, was used, since it provides a high-level interface for exploring complex datasets. This library enables effective visualization, interpretation, and comparison of the results obtained in the research.
@@ V-A Scenario 1: Static Positions with Beam-Steering Optimization
TABLE III: Study of GA parameterisation.
@@
Figure 7 shows the evolution of the OF value for a beam-steering configuration, as well as the constant baseline value corresponding to the use of omnidirectional communication. The figure shows that the application of beam-steering antennas can effectively improve the performance of communications: the best fitness value for the omnidirectional configuration is 0.11, in contrast with the value obtained for the beam-steering configuration, which is approximately 350000.
A specific example is depicted in Figure 8, which was obtained after a computation time of 29.3 s. This result corresponds to an average data rate of 2856 bps and a bottleneck data rate of 129 bps. In contrast, for the omnidirectional configuration, the average data rate was 0.89 bps and the bottleneck data rate was 0.12 bps. Figure 8 shows the direction of the antenna taken by each UAV. It is easy to notice that the beam-steering solution favors pointing the beams away from the jammer (e.g., UAV0 and UAV3), and/or towards other UAVs (e.g., UAV1 and UAV2).
@@ V-B Scenario 2: Fixed Swarm Deployment Area with Beam-Steering and UAV Position Optimization
Figure 10: Scenario 2: convergence of the OF in acquiring the positions.
@@
In the omnidirectional communication configuration, GA can obtain a very fine-tuned arrangement of the UAVs in a significantly shorter time interval, trying to place the swarm as far as possible from the jammer. Its results are shown in Figure 12. Although the positioning of the UAVs using the beam-steering antenna configuration are obviously not ideal when compared to the omnidirectional configuration, there is a drastic decrease in communication performance when using omnidirectional antennas.
@@ V-C Scenario 3: Mobile Swarm Deployment Area
In this dynamic scenario, the UAV swarm continuously moves in a straight line towards some mission objective, while the UAVs can change their relative positions within its deployment area, as well adapt the beam-steering configuration accordingly. Figure 13 illustrates the scenario. Figure 14 shows the system execution flowchart. Additional simulation parameters for Scenario 3 are listed in Table V.
@@
The initial UAV swarm situation is depicted in Figure 15.
@@
Selected snapshots of the UAV swarm movement are depicted in Figure 16. Despite processing time constraints, the GA could effectively cluster UAVs and direct antennas to mitigate jamming effects. Over the simulation, the sequence of objective function curves (omitted due to page limits) reflect a steady optimization of communication quality, as seen in evolving UAV topologies as the swarm operational area slides horizontally. The success in aligning antennas and adjusting positions demonstrates the GA capability to provide robust anti-jamming strategies within the given constraints, even if the found solutions were still not optimal. It must be noted that the GA seeks optimal UAV positions and beam-steering solutions for specific time instants of the swarm movement. Although the GA could improve its decisions for longer time windows, the latter would also mean that the swarm would adapt its topology and beam-steering less frequently. For very distant jammers, enlarging the time window is possible, since angular directions change slowly, otherwise, enlarging the time window will also lead to non-optimality. Optimization of the time window length is left for future work.